OpenStudy (anonymous):

Which is less than zero point three five? three-tenths eleven twenty-fifths two-fifths nine-twentieths

OpenStudy (diyadiya):

three-tenths = 3/10 =0.3 eleven twenty-fifths= 11/25=0.44 two-fifths = 2/5=0.4 nine-twentieths=9/20=0.45 Now can you tell me which one is less than 0.35?

OpenStudy (anonymous):

0.3

OpenStudy (diyadiya):

Right! :) 0.3 = 3/10 = three tenths :)

OpenStudy (anonymous):

thank you

OpenStudy (diyadiya):

Welcome! :)